OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x – y = 7 x = y + 3 A. x – y = y + 3 B. (y + 3) – y = 7 C. x – (y + 3) = 7 D. x = (7 – x) + 3

OpenStudy (anonymous):

sub x =y+3 into x-y=7 (y+3)-y=7

OpenStudy (igreen):

Substitute y + 3 for x in the first equation: x - y = 7 Replace x with y + 3, what do you get? @fangs

OpenStudy (anonymous):

thank you for the help

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @fangs

OpenStudy (igreen):

@Let.Van Don't give out answers.

OpenStudy (anonymous):

is a direct substitution...The best way to learn it is to see as many of these as possible